OpenStudy (anonymous):

find f(x) and g(x) so that the function can be described as y = f(g(x)). y =9/x^2 + 2

OpenStudy (mathmale):

You are given a "composite function here. This means that one function is the INPUT of the other function. Have you done this kind of problem before?

OpenStudy (mathmale):

Note that x^2 could be regarded as a separate function.

OpenStudy (pawanyadav):

This question should contains options.

OpenStudy (pawanyadav):

Because we have more than one choice.

OpenStudy (mathmale):

I agree that there may be more than just one solution. Review: a composite function is one in which one function g(x) is the input to another function f(x). Try again. Do you see a separate function in y =9/x^2 + 2?

OpenStudy (danjs):

let g(x) be some part of what the right side is f(that part) = y the whole function

OpenStudy (danjs):

g(x) = x f(x) = 9/g(x)^2 + 2 f(g(x)) = 9/x^2 + 2 or g(x) = x^2 f(x) = 9/g(x) +2 f(g(x)) = 9/g(x) + 2 or g(x) = g(x) = 9/x^2 f(x) = g(x) + 2